OpenStudy (anonymous):

What value is a discontinuity of x^2+2x+3/x^2-x-12? x = -1 x = -2 x = -3 x = -4

OpenStudy (cwrw238):

this will occur when the denominator x^2 - x - 2 = 0

OpenStudy (anonymous):

ok

OpenStudy (cwrw238):

can you solve x^2 - x - 2 = 0 for x?

OpenStudy (anonymous):

yes using the quadratic formula correct?

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

would it be one @cwrw238

OpenStudy (cwrw238):

no you can factor this one x^2 - x - 2 = x^2 - 2x + x - 2 = x(x - 2) + 1(x - 2) = (x - 2)(x + 1) so x = -1 , 2

OpenStudy (cwrw238):

so its -1 in your list